Milhouses demand function for flood pants is represnted by the equation Qd=20-(p/2). The price of flood pants rises from 10 to 14. What is milhouses price elasticity of demand between these two points?
Answer #1

Answer

Qd=20-(p/2)

p=10 then Qd=20-(10/2)=15

p=14 then Qd=20-(14/2)=13

Elasticity of demand=(change in quantity/average quantity)/(change in price/average price)
Change in quantity=13-15=-2
average quantity=(13+15)/2=14
change in price=14-10=4
average price=(14+10)/2=12
Elasticity of demand=(-2/14)/(4/12)

=-0.428571429

=-0.43

=0.43 ( absolute value)

the milhouses price elasticity of demand between these two points is 0.43 and it is inelastic as the elastcity is below 1.